a tennis ball has a mass of 0.06kg and has kinetic energy of 2.94KJ when served. what is it's speed?


1
Expert's answer
2021-03-08T08:30:42-0500

By defintion, the kinetic energy is given as follows:


"K = \dfrac{mv^2}{2}"

where "m = 0.06kg" is the mass, and "v" is the speed of the ball. Expressing "v" from this formula and substituting "K = 2.94\times 10^3J", obtain:


"v = \sqrt{\dfrac{2K}{m}}\\
v = \sqrt{\dfrac{2\cdot 2.94\times 10^3}{0.06}}\approx 313m/s"

Answer. 313 m/s.